OpenStudy (anonymous):

A deposit of $5,000 is made into a savings account that offers 7.5% annual interest. Which equation models the amount of money in the account after t years? P(t) = 5,000(1.75)t P(t) = 5,000(1.075)t P(t) = 5,000(0.925)t P(t) = 5,000(0.25)t

OpenStudy (anonymous):

P(t) = A(1 + r)^t WHERE A = Amount r = decimal rate IE 80% = .8 so 1+r = 1.8 t = term in years the next thing you'll want to know is what about x terms per year, like monthly. in that case P=A(1 + r/n)^(nt) WHERE n= periods per term From there you can work out just about any of these problems